OpenStudy (anonymous):

I've tried working this out, but it isn't working out for me for any of these.

OpenStudy (anonymous):

None of these equal each other when plugged in. What am I doing wrong?

OpenStudy (anonymous):

3x+15=6x-9 15+9 = 6x - 3x 21 = 3x x = 21/3 = 7

OpenStudy (anonymous):

What are you saying the answer is?

OpenStudy (anonymous):

sorry I didn't see the bottom, but it's the last one 8.

OpenStudy (anonymous):

try subbing it in.

OpenStudy (anonymous):

Oh, ok. I was making a slight calculation error. I got what I was doing wrong. Stupid mistakes. Thanks!

OpenStudy (anonymous):

no problem :)